OpenStudy (anonymous):

EVERYONE WHO ANSWERS RIGHT GETS MEDALS!!!! If a pendulum with length 169 cm has a frequency of 1/20 Hz, find the length of a pendulum that has frequency 1/5 Hz, given that you know that the frequency varies inversely with the square root of the length.

jhonyy9 (jhonyy9):

169 ---- 1/20 Hz x ----- 1/5 Hz ----------------- x=(169 x 1/5)/1/20 = (169 x 20)/5 =169 x 4 = 42,2 cm

OpenStudy (anonymous):

I think using the formula of simple pendulum T = 2pi * squrt(L/g) to solve for g 20 = 2pi * squrt(1.69/g) => g= 1.69 * (2pi/20)^2=0.1668 m/s^2 then 5 = 2pi* squrt(L/0.1668) => L= (5/2pi)^2*0.1668 = 0.106 m = 10.6 cm